OpenStudy (anonymous):

How do you simplify 6sin5xcos5x?

OpenStudy (anonymous):

3sin(10x)

OpenStudy (raden):

6sin5xcos5x = 3 * 2sin5xcos5x = 3 sin(2*5x) = 3sin(10x)